### v1

# CPRet: A Dataset, Benchmark, and Model for Retrieval in Competitive Programming

Abstract
Competive programming benchmarks are widely used in scenarios such as programming contests and large language model assessments. However, the growing presence of duplicate or highly similar problems raises concerns not only about competition fairness, but also about the validity of competitive programming as a benchmark for model evaluation. In this paper, we propose a new problem—similar question retrieval—to tackle the problem. Due to the lack of both data and models, solving this problem is challenging. To this end, we introduce CPRet, a retrieval-oriented benchmark suite for competitive programming, covering four retrieval tasks: two code-centric (i.e., Text-to-Code, Code-to-Code) and two newly proposed problem-centric tasks (i.e., Problem-to-Duplicate, Simplified-to-Full)—built from a combination of automatically crawled problem–solution data and manually curated annotations. Our contribution includes both high-quality training data and temporally separated test sets for reliable evaluation. Besides, we further develop two task-specialized retrievers based on this dataset: CPRetriever-Code, trained with a novel Group-InfoNCE loss for problem–code alignment, and CPRetriever-Prob, fine-tuned for indentifying problem-level similarity. Both models achieve strong results and are open-sourced for local use. Finally, we analyze LiveCodeBench and find that high-similarity problems inflate model pass rates and reduce differentiation, underscoring the need for similarity-aware evaluation in future benchmarks.
1 Introduction
Competitive programming contests—from high-school Olympiads like the IOI to university-level events such as the ICPC—challenge participants to solve algorithmic problems under tight time and memory limits, requiring both strong coding skills and deep algorithmic insight. Because problem statements are precisely specified in natural language and solutions can be graded automatically, these problems have become a canonical benchmark for assessing the reasoning and coding abilities of large language models (LLMs) livecodebench ; humaneval . Recent advances in code-oriented LLMs—spanning commercial flagships such as OpenAI o4-mini o3_o4mini , Gemini-2.5-Pro gemini25pro , and Grok-3-Mini (High) grok3 , alongside research releases like DeepSeek-R1 deepseek-r1 highlight the value of competitive programming as a testbed for algorithmic reasoning, program synthesis, and computer-science education.
Competitive programming has witnessed rapid growth over the past three decades, with thousands of new problems introduced annually. This expansion has led to a growing concern: the increasing presence of similar or repetitive problems within large repositories. As shown in Figure 1, our collected data reveals a significant rise in community discussions around duplicate problems in recent years. However, it is often difficult to determine whether a new problem is a duplicate of existing ones, as this typically relies on the memory and judgment of human problem setters. Unchecked duplication brings concrete downsides. In human programming contests, repeated or highly similar problems give an unfair advantage to participants who have seen them before. In competitive programming benchmarks for LLMs, the presence of repeated or highly similar problems can lead to inflated performance scores, as models may rely on memorized instances of past competition problems from their training data rather than demonstrating genuine algorithmic reasoning. This compromises the evaluation of their true ability to solve novel and unfamiliar challenges.
Figure 1: Year-wise trend of annotated duplicate problem discussions.
A promising way to address the issue of duplicate problems is to leverage retrieval models that identify semantically similar problems based on learned representations. Such models offer a scalable and model-agnostic solution for redundancy detection in large problem repositories. However, despite their potential, there is a lack of established benchmarks specifically targeting problem-level retrieval. Prior efforts in competitive programming retrieval have primarily focused on code-centric tasks, such as retrieving solution code given a problem description (Text-to-Code) or retrieving alternative correct solutions given a reference implementation (Code-to-Code), which overlook the dimension of retrieving similar problems themselves—an essential capability for both redundancy detection and educational search.
To fill this gap, we introduce two new problem-centric retrieval tasks. (I) Problem-to-Duplicate asks a model to retrieve the most semantically or structurally similar problem to a given target. We manually annotated 700 duplicate pairs; 200 pairs form a held-out test set in which the model must locate the duplicate among a corpus of 10,000+ problems, and the remaining $\sim 500$ pairs are used for training. (II) Simplified-to-Full supplies a simplified or paraphrased statement and requires the model to find the original full version. We collected 17,000 human-written simplifications aligned with their source problems; 10,000 pairs constitute the test set and the remaining $\sim 7,000$ pairs serve as training data. These tasks target practical problem-retrieval needs—spotting duplicate or highly similar problems and retrieving the original statement from a simplified description. In addition, we perform a temporal analysis of the two code-centric tasks and uncover severe train-test leakage: models fare markedly worse on the newest problems. To this end, we rebuild both benchmarks, drawing the test sets exclusively from the most recent problems to enforce strict temporal separation.
Furthermore, to address the real-world retrieval demands, we develop two task-specialized models CPRetriever-Code and CPRetriever-Prob using our new corpus: (I) the first one is trained on 38 k problems and 2.9 M multi-language solutions with a new Group-InfoNCE loss that treats every correct solution of the same problem as a positive and penalizes similarity within that positive set. This encourages alignment across diverse implementations, and it achieves the overall best performance on all code-centric tasks. (II) The second one is fine-tuned from CPRetriever-Code on our duplicate-detection and simplified-to-full datasets, and it performs the overall best results on the two problem-centric tasks.
Finally, we apply CPRetriever-Prob to analyze problems in LiveCodeBench released after September 2024 by computing each problem’s maximum similarity to earlier data. We observe two key trends: (I) Pass rates increase with similarity—problems more similar to prior examples are consistently easier for models to solve, regardless of difficulty; (II) High similarity reduces model differentiation—as similarity grows, performance gaps between models shrink, suggesting that models may solve such problems by relying on memorized patterns rather than demonstrating true reasoning ability in novel scenarios. These findings highlight the importance of accounting for problem similarity when constructing benchmarks to avoid overestimating model capabilities.
Overall, our key contributions can be summarized as follows:
$\bullet$ Comprehensive Benchmark: We construct a large-scale, high-quality dataset for competitive programming, covering four key retrieval tasks (two focused on code retrieval and two newly proposed for problem retrieval). Our benchmark reflects realistic retrieval needs in programming contests and addresses issues such as problem repetition and dataset leakage.
$\bullet$ Effective Models: We develop two task-specialized retrieval models with strong performance, trained with our newly designed Group-InfoNCE loss. Both models are open-sourced and can be run locally—an important feature given the privacy constraints in contest settings—providing practical tools for code and problem retrieval and helping mitigate the growing issue of duplicate problems in competitive programming.
$\bullet$ Similarity-Aware Evaluation: We provide insightful studies on the impact of problem similarity in competitive programming, and find that problems with high similarity to prior data yield inflated pass rates and reduced model differentiation, especially for weaker models. These results highlight the importance of controlling for similarity when designing benchmarks and suggest that future datasets should stratify or filter test problems based on their similarity to past content.
2 Related Work
2.1 Datasets for Competitive Programming
Several datasets have been proposed to train and evaluate LLMs on competitive programming tasks. Early efforts include Description2Code Description2Code , which gathered problems and solutions from CodeChef and Codeforces, followed by the APPS dataset apps , which became a widely used benchmark across a mix of competition-style and introductory problems. CodeContests alphacode , used in DeepMind’s AlphaCode, expanded coverage but remained limited in scale. More recently, TACO TACO constructed the largest dataset (26K problems) by aggregating multiple sources. In the retrieval setting, CoIR CoIR built on APPS and CodeTransOcean codetransocean to define two retrieval tasks: text-to-code and code-to-code, helping assess model capabilities in aligning problem descriptions and solution code. Despite their contributions, these datasets have several limitations: (i) many were collected before 2023 and suffer from potential data leakage, (ii) problem types are often narrow in scope, and (iii) both problem statements and solutions are typically limited to English and Python, reducing linguistic and implementation diversity.
For evaluating programming ability, HumanEval humaneval remains a widely used standard, featuring hand-written Python problems with corresponding unit tests for functional correctness. LiveCodeBench livecodebench improves upon earlier efforts by continuously collecting real-world problems from LeetCode, AtCoder, and Codeforces, and applies temporal separation to reduce data contamination. However, these benchmarks do not account for problem-level similarity between new and historical problems—a distinct issue from data leakage—which can still lead to overly optimistic assessments of model performance due to hidden redundancy.
2.2 Embedding Models for Retrieval
Dense retrieval models learn to map queries and candidates into a shared embedding space for efficient similarity search. Early methods like DSSM dssm and Sentence-BERT sentencebert paved the way for supervised contrastive training, leading to influential retrievers such as Dense Passage Retrieval (DPR) dpr .
Recent efforts have focused on building general-purpose embedding models (e.g., E5-Mistral e5mistral , GTE-Qwen2 gteqwen2 , SFR-Embedding sfrtext , NV-Embed nvembed ; nv-retriever , Linq-EmbedLinq ) that perform well across a wide range of retrieval tasks. These models are commonly evaluated on MTEB MTEB , a standardized benchmark suite covering diverse tasks such as semantic search, classification, and reranking. For the code domain, specialized models like CodeSage codesage , SFR-Embedding-Code sfrcode , Qodo-Embedqodo use code-specific data and retrieval-aware objectives to significantly improve performance on benchmarks such as CoIR. While both general-purpose and code-specific embedding models can be applied to competitive programming, we find their performance on code-centric and problem-centric retrieval tasks remains limited—highlighting the need for a dedicated solution tailored to the unique characteristics of this domain.
3 Method
3.1 Overview of Retrieval Tasks
We define and evaluate four distinct retrieval tasks, each capturing a different dimension of semantic understanding and reuse in competitive programming:
  • Text-to-Code Retrieval: Given the full natural language description of a problem, retrieve one or more correct solution codes. This task assesses a model’s ability to align problem semantics with executable implementations.
  • Code-to-Code Retrieval: Given one accepted solution, retrieve alternative correct solutions to the same problem. This task evaluates a model’s capacity to understand code functionality and identify semantically equivalent but syntactically diverse implementations.
  • Problem-to-Duplicate Retrieval: Given a problem description, retrieve other problems that are duplicated, including exact matches or closely related in terms of solution strategy. This task is useful for identifying redundancy and measuring problem novelty across platforms.
  • Simplified-to-Full Retrieval: Given a simplified version of a problem, retrieve the corresponding one with full description. This task examines cross-abstraction retrieval capabilities, bridging accessibility-oriented rewrites and original problem statements.
These tasks jointly form a comprehensive benchmark for evaluating retrieval models in the context of competitive programming, covering problem-code alignment, solution diversity, duplication detection, and abstraction-level matching. Table 1 summarizes the statistics of each retrieval task, including training and test set sizes, as well as average token lengths for both queries and corpus items.
Table 1: Statistics of the four retrieval tasks in CPRet, including the number of training and test items. $L_{\text{Query}}$ and $L_{\text{Corpus}}$ denote average token lengths.
Retrieval Task #Train- Problem #Train- Code #Train- Pair #Test- Query #Test- Corpus #Test- Qrels $L_{\text{Query}}$ $L_{\text{Corpus}}$
Text-to-Code 38.8K 2.93M 2.93M 4.9k 41.6k 41.6k 1038 1210
Code-to-Code 4.8k 39.8k 39.8k 1132 1185
Problem-to-Duplicate 874 / 491 168 10.9k 202 565 765
Simplified-to-Full 7.6k / 7.6k 10k 10k 10k 226 697
3.2 Dataset and Benchmark Construction
3.2.1 Problems and Codes Data
We construct a large-scale dataset of programming contest problems and their corresponding accepted solutions to support both model training and the evaluation of retrieval tasks. Compared to existing datasets, our dataset provides broader temporal coverage, richer language diversity, and more varied contest formats. It includes recent problems from multiple online judges, spans several programming languages, and covers both ICPC-style (International Collegiate Programming Contest) and OI-style (Olympiad in Informatics) problems—the latter involving partial scoring and more complex algorithmic requirements. Each problem is paired with one or more solutions and annotated with precise timestamp information, enabling temporally-aware training and evaluation across semantic retrieval tasks. As summarized in Table 2, our dataset—CPRet-PCPCD—offers wider coverage across problem sources, types, and languages, and serves as a stronger foundation for retrieval-based modeling in competitive programming.
Table 2: Comparison of competitive programming datasets. #Src: Number of data sources (e.g., online judges). #P-Type: Number of problem types (e.g., ICPC-style with full-score only vs OI-style with partial scoring). #Lang-P: Number of languages used for problem descriptions. #Lang-C: Number of programming languages used in solutions. See Appendix A.3 for more detailed dataset statistics.
Dataset #Prob #Code #Src #P-Type #Lang-P #Lang-C Cut-off
Description2Code Description2Code 7.8K 309K 3 1 1 2 2016/08
APPSapps 10K 232K 7 1 1 1 2020/10
CodeContestsalphacode 13.6K 4.5M 5 1 1 4+ 2021/07
TACOTACO 26.4K 1.55M 10 1 1 1 2023/02
CPRet-PCPCD (ours) 42.2K 2.9M 12 2 3 20+ 2024/12
As shown in Figure 2, on the Text-to-Code task using the historical APPS dataset, model performance improves over time as newer and stronger models are introduced. However, we observe a clear drop in performance when problems are grouped by their original release year—only stabilizing after 2022. This indicates potential data leakage or memorization in earlier benchmarks and underscores the need for temporally separated evaluation. Accordingly, we use problems and solutions from 2023 onward in our collected dataset as the test set for both the Text-to-Code and Code-to-Code tasks.
Figure 2: Left: NDCG@10 scores on CoIR/APPS across model release dates, showing steady performance improvement as models evolve. Right: Performance on the Text-to-Code task grouped by problem release year, revealing sharp degradation on older problems, especially around major dataset release points (Description2Code, APPS, CodeContests).
3.2.2 Duplicate Programming Problem Pairs
In many instances, when a newly released contest problem closely resembles or duplicates a previously published one, participants will highlight this in the associated discussion threads or contest forums. To leverage this observation, we collected all publicly available discussion threads and blog posts from two major competitive programming platforms: Codeforces and Luogu. We applied a combination of keyword-based heuristics and large language model (LLM)-based classification to identify approximately 5,000 potentially relevant entries. These candidates were subsequently verified through manual annotation by several experienced competitive programmers.
To ensure consistency and clarity in annotation, we defined three levels of duplication between problems:
  • Exact Match: An accepted (AC) solution for one problem can directly pass the other without modification.
  • Near Match: An AC solution for one problem can be adapted to solve the other with minor edits that could reasonably be made by someone unfamiliar with the second problem.
  • Method Match: The core idea and solution approach are the same, but the code differs in non-essential or implementation-specific details.
We ultimately identified around 700 pairs of duplicate problems. Since some problems may belong to duplication clusters involving more than two problems, we first performed clustering to group mutually duplicate problems. We then randomly selected 30% of the clusters to construct the test set for the Problem-to-Duplicate retrieval task. Within each selected cluster, one problem was randomly designated as the query, while the remaining ones were placed in the corpus. To further increase task difficulty and realism, we added all other Codeforces problems to the corpus as distractors.
3.2.3 Simplified and Full Problem Description Pairs
To support beginners and non-native speakers, users on the Luogu platform have contributed Chinese translations and simplified versions of competitive programming problems originally written in English (e.g., on Codeforces) or Japanese (e.g., on AtCoder). We crawled these user-generated simplifications and applied filtering procedures to remove low-quality entries, particularly those stemming from direct or unedited machine translation.
After cleaning, we obtained approximately 17,000 high-quality pairs of full problem descriptions and their corresponding simplified versions. From this set, we randomly selected 10,000 pairs to construct the test set for the Simplified-to-Full retrieval task, where the simplified description serves as the query and the full version is used as the corpus. The remaining around 7,000 pairs are used for training.
3.3 Competitive Programming Retrievers
3.3.1 CPRetriever-Code: Multi-Positive Contrastive Learning for Problem-Code Alignment
We train CPRetriever-Code using a supervised contrastive learning framework tailored for aligning problem descriptions with diverse correct solutions. Each problem is paired with multiple valid codes, and the model learns to capture this alignment while preserving representation diversity.
We begin with the standard InfoNCE loss oord2018representation , where each (problem, solution) pair $(x_{i},x_{i}^{+})$ is contrasted against others in the batch:
$\mathcal{L}_{\text{InfoNCE}}=-\log\frac{\exp(\mathrm{sim}(x_{i},x_{i}^{+})/\tau)}{\sum_{j=1}^{N}\exp(\mathrm{sim}(x_{i},x_{j}^{+})/\tau)}.$ (1)
Limitation of Single-Positive Loss. In practice, many problems have multiple correct solutions $G_{i}={x_{i}^{1+},\ldots,x_{i}^{m+}}$. A common extension averages the similarities over all positives zhao2023leveraging :
$\mathcal{L}_{\text{MultiPos}}=-\log\frac{\sum_{k=1}^{m}\exp\left(\mathrm{sim}(x_{i},x_{i}^{k+})/\tau\right)}{\sum_{j\neq i}^{N}\exp\left(\mathrm{sim}(x_{i},x_{j})/\tau\right)}.$ (2)
However, this ignores the internal structure of the positive set and does not encourage consistency among the positives.
Group-InfoNCE. To better utilize multiple correct solutions, we propose Group-InfoNCE, which treats the positive set as a whole. The loss promotes similarity between the query $x_{i}$ and its group $G_{i}$, while contrasting against other problems and groups:
$\mathcal{L}_{\text{Group}}=$ $-\log\frac{\exp\left(\mathrm{sim}_{G}(x_{i},G_{i})/\tau\right)}{\exp\left(\mathrm{sim}_{G}(x_{i},G_{i})/\tau\right)+\sum_{j\neq i}\left[\exp\left(\mathrm{sim}(x_{i},x_{j})/\tau\right)+\exp\left(\mathrm{sim}_{G}(x_{i},G_{j})/\tau\right)\right]}$ (3)
$+\frac{\text{Penalty}_{G}(x_{i},G_{i})}{\tau^{2}}.$
Here, group similarity is defined as:
$\mathrm{sim}_{G}(x_{i},G_{j})=\frac{1}{m}\sum_{k=1}^{m}\mathrm{sim}(x_{i},x_{j}^{k+}),$ (4)
and the variance-based regularization encourages consistency within the group:
$\text{Penalty}_{G}(x_{i},G_{i})=\mathrm{Var}_{k=1}^{m}\left(\mathrm{sim}(x_{i},x_{i}^{k+})\right).$ (5)
This group-based formulation improves representation quality by explicitly modeling the structure of multiple correct outputs, leading to more stable and discriminative embeddings for code-centric tasks.
Format Masking for Robustness. To reduce reliance on superficial cues, we apply random masking to non-essential parts of the problem description during training, including I/O format explanations, sample inputs/outputs, and explicit data constraints. This encourages the model to focus on the core algorithmic intent rather than overfitting to formatting patterns, and improves generalization to problems with diverse or unfamiliar layouts.
3.3.2 CPRetriever-Prob: Fine-Tuning on Problem-Level Tasks
CPRetriever-Prob is obtained by fine-tuning CPRetriever-Code on problem-level retrieval tasks. While CPRetriever-Code captures general alignment between problems and solutions, this stage specializes the model for retrieving semantically or structurally related problems—such as duplicates or simplified versions.
In particular, we jointly train on data from the Problem-to-Duplicate and Simplified-to-Full tasks (Section 3.1), enabling the model to learn fine-grained alignment between problems that are either functionally equivalent or differ primarily in abstraction level.
Training Format and Loss. Each training example is structured as a triplet $(x,x^{+},x^{-})$, where $x$ is the query, $x^{+}$ is a semantically aligned problem, and $x^{-}$ is a hard negative. We use the triplet margin loss schroff2015facenet :
$\mathcal{L}_{\text{triplet}}=\max\left(0,\mathrm{sim}(x,x^{-})-\mathrm{sim}(x,x^{+})+\alpha\right),$ (6)
where $\mathrm{sim}(\cdot,\cdot)$ is cosine similarity and $\alpha$ is the margin.
Hard Negative Mining. As our dataset primarily contains positive pairs, we construct hard negatives by first retrieving the top-10 most similar candidates for each query using a pretrained retriever (SFR-Embedding-Code-2B_Rsfrcode ) and randomly sampling one as a negative. To ensure the sampled negative is not a false match, we use Qwen-2.5-Maxqwen2 to verify that it is not equivalent to the query in terms of problem intent. This automatic filtering yields challenging yet reliable negatives without requiring manual annotation.
4 Experiment
4.1 Experimental Setup
We develop two retrieval models based on the SFR-Embedding-Code-2B_Rsfrcode embedding model, which is built on top of Gemma-2-2Bgemma and further pre-trained on code-related tasks. Detailed training configurations are provided in Appendix A.1.
4.2 Main Results
Table 3: Retrieval performance across models with different scales. We use NDCG@10 jarvelin2002cumulated as the primary metric to measure ranking quality across tasks. T2C: Text-to-Code. C2C: Code-to-Code. P2Dup: Problem-to-Duplicate. S2Full: Simplified-to-Full. Avg: Average of the four tasks.
Scale Model Type Size T2C C2C P2Dup S2Full Avg
Tiny contriever-msmarco contriever-msmarco general 109M 1.06 7.49 5.26 43.04 14.21
multilingual-e5-small me5 general 110M 2.73 13.52 14.90 52.07 20.80
codesage-small-v2 codesage code 130M 10.16 22.24 17.42 67.07 29.22
gte-modernbert-base mgte general 149M 14.99 36.22 21.12 77.45 37.44
Small bge-large-zh-v1.5 bge general 324M 2.58 7.49 14.75 37.34 15.54
stella_en_400M_v5 stella general 400M 2.12 9.76 13.14 57.56 20.64
multilingual-e5-base me5 general 278M 1.96 13.66 17.24 58.26 22.78
bge-m3 bge-m3 general 569M 5.69 12.67 20.12 63.81 25.57
codesage-base-v2 codesage code 356M 17.30 29.38 19.97 74.36 35.25
SFR-Emb.-Code-400M sfrcode code 400M 9.43 43.59 19.40 75.31 36.93
Medium multilingual-e5-large me5 general 560M 4.27 18.51 19.19 65.02 26.75
multi.-e5-large-instruct me5 general 560M 6.64 28.84 23.31 61.28 30.02
gte-Qwen2-1.5B-inst. gte general 1.5B 10.76 23.41 27.06 69.15 32.60
stella_en_1.5B_v5 stella general 1.5B 9.22 21.40 29.45 72.91 33.24
inf-retriever-v1-1.5b inf general 1.5B 18.31 28.17 30.11 74.19 37.70
codesage-large-v2 codesage code 1.3B 20.78 35.23 22.43 78.70 39.28
Qodo-Embed-1-1.5B qodo code 1.5B 22.93 36.52 33.37 84.05 44.22
SFR-Embedding-2 sfrtext general 2B 20.57 50.38 35.96 73.02 44.98
SFR-Emb.-Code-2B sfrcode code 2B 39.60 68.05 45.26 86.43 59.84
CPRetreiver-Code code 2B 70.40 70.59 38.68 81.45 65.28
CPRetreiver-Prob code 2B 56.50 70.68 60.06 90.74 69.50
Large GritLM-7B Grit general 7B 0.22 8.74 11.18 29.81 12.49
NV-Embed-v2 nv-retriever ; nvembed general 7B 7.09 34.88 25.49 63.14 32.65
inf-retriever-v1 inf general 7B 17.43 25.38 30.85 78.46 38.03
gte-Qwen2-7B-instruct gte general 7B 17.96 30.72 35.95 78.41 40.76
SFR-Emb.-Mistral sfrtext general 7B 22.15 50.92 31.88 69.38 43.58
Linq-Embed-Mistral Linq general 7B 21.99 52.06 36.51 72.79 45.84
Qodo-Embed-1-7B qodo code 7B 36.47 51.91 47.15 91.17 56.68
We evaluate over 20 strong embedding models drawn from top-performing entries on the MTEBMTEB and CoIRCoIR benchmarks, as summarized in Table LABEL:tab:model_comparison. Models trained specifically for the code domain continue to outperform general-purpose models by a significant margin. Among them, our two proposed models—CPRetriever-Code and CPRetriever-Prob—achieve the best overall results across tasks.
CPRetriever-Code, trained with Group-InfoNCE on problem-code pairs, excels on the code-centric tasks—Text-to-Code (T2C) and Code-to-Code (C2C). After fine-tuning on problem-level data, CPRetriever-Prob achieves large gains on the problem-centric tasks—Problem-to-Duplicate (P2Dup) and Simplified-to-Full (S2Full)—with a modest drop on C2C and a more noticeable decline on T2C.
We attribute the T2C degradation to differing retrieval demands: T2C relies on implementation-specific details in problem descriptions, whereas P2Dup and S2Full emphasize higher-level semantic similarity. This trade-off is also evident in other models—for instance, Qodo-Embed-1-7Bqodo achieves the best S2Full score but underperforms on T2C compared to SFR-Embedding-Code-2B_Rsfrcode . These observations motivate our decision to release two task-specialized models, each optimized for different objectives. Further analysis of this trade-off is provided in Appendix A.2.2, and detailed ablation studies on the Group-InfoNCE loss appear in Appendix A.2.1.
4.3 Impact of Problem Similarity on Model Success in Competitive Programming Tasks
To assess how prior problem similarity influences model performance, we examine 388 LiveCodeBench problems released after September 1, 2024. For each, we compute the average pass rate across all evaluated models and measure its maximum cosine similarity to earlier CPRet-PCPCD problems using CPRetriever-Prob.
As shown in Figure 3, we observe a clear trend: on average, problems with higher maximum similarity to past data have significantly higher pass rates. This pattern holds across difficulty levels, but with notable differences: (I) Medium-difficulty problems show the strongest positive correlation, suggesting that models benefit substantially from prior exposure to semantically similar content. (II) Hard problems exhibit generally low pass rates, though a slight upward trend remains visible with increasing similarity. (III) Easy problems maintain consistently high pass rates overall. Due to a few low-performing outliers at high similarity, the regression line shows a slight downward slope—likely an artifact of skewed distribution rather than a true negative effect. The binned box plot (right) further confirms this relationship, reinforcing that semantic similarity to known problems is a key factor in model success.
Figure 3: Impact of similarity to prior problems on pass rate. Left: For 388 post-2024.9.1 LiveCodeBench problems, we plot average model pass rate vs. maximum similarity to historical CPRet problems, computed via CPRetriever-Prob. Regression lines are shown by difficulty, with the gray line indicating the overall trend across all problems. Right: Pass rates increase monotonically across similarity bins, confirming a strong link between retrieval similarity and generation success.
To examine how model performance relates to problem similarity, we analyze evaluation results from OpenAI’s O3-Mini and O4-Mini variants as reported in LiveCodeBench. Each model is available in three modes—Low, Medium, and High—which correspond to increasing levels of resource consumption (e.g., longer context length, higher inference cost, and possibly larger internal activations). As shown in the left panel of Figure 4, we observe that the performance gap between these variants narrows as the maximum similarity to past problems increases, and nearly vanishes in the highest bin (0.80–0.90). In contrast, on low-similarity problems, higher-capacity models (especially the High variants) maintain strong pass rates, while lower-capacity models degrade significantly. This suggests that high-similarity problems may allow weaker models to perform well by potentially leveraging memorized patterns or surface-level matching, while low-similarity problems are more likely to expose differences in model capability—such as generalization and robustness in reasoning. The right panel shows that easy problems have slightly higher similarity to past data than hard ones, but the difference is modest. This indicates that similarity is not determined by difficulty alone, and should be treated as an independent factor when constructing benchmarks.
Figure 4: Left: Models with higher capacity (High) maintain relatively high performance even on low-similarity problems, while lower-capacity models benefit more from high-similarity problems. Right: Easy problems show slightly higher maximum similarity to past problems than hard ones, but the difference is limited.
5 Limitations and Future Work
Ongoing test set leakage. Although we use the most recent problems (from 2023–2024) to construct temporally-separated test sets for code-centric tasks, these examples are likely to appear in future model training corpora as web-scale datasets are updated. As a result, benchmark performance may become inflated over time, even without intentional misuse. To address this, we plan to release regular updates—every 6 to 12 months—to maintain a clean test split while ensuring the retrieval corpus remains diverse and representative.
Incomplete and emerging duplicate problems. Despite 700 annotated pairs in our Problem-to-Duplicate task, many duplicates—past and future—remain undiscovered, especially on decentralized or low-resource platforms. To address this, we are developing an open-source retrieval platform that supports community-driven annotations and helps contest organizers detect redundancy pre-publication.
Limited analysis beyond programming. While our study focuses on competitive programming, duplication and high similarity may also affect evaluation in other domains, such as mathematical competitions (e.g., IMO, AIMO). These issues remain underexplored despite their potential to undermine fair assessment. We encourage future work to examine problem reuse and redundancy in such benchmarks.
Appendix A Technical Appendices and Supplementary Material
A.1 Additional Experimental Setup Details
For our contrastive pretraining, we use a batch size of 1024, enabled by gradient caching[41] to optimize GPU memory usage. The model is trained with a learning rate of $3\times 10^{-8}$, temperature $\tau=0.07$, and weight decay of 0.01, using the AdamW[42] optimizer. Each query is paired with $m=16$ positive samples, randomly sampled from the dataset. If fewer than $m$ positive samples are available, duplicates are used to pad the group. The model is trained for 20 epochs on 8 NVIDIA A800 GPUs, taking approximately 70 hours to complete. This stage focuses primarily on problem-code representation alignment. The resulting model, which demonstrates strong performance on code-related retrieval tasks, is denoted as CP-Retriever-Code.
To improve the model’s performance on problem-problem retrieval tasks, we perform a second-stage fine-tuning using a multi-task setup. We combine training data from the Problem-to-Duplicate and Simplified-to-Full tasks, downsampling each to approximately 1,000 examples to ensure balanced representation. The two subsets are then mixed to form the final fine-tuning dataset. This stage uses a batch size of 1, learning rate of $2\times 10^{-6}$, and weight decay of 0.01. The model is trained for 1 epoch, which takes approximately 1 hour on a single NVIDIA A800 GPU. While this model, referred to as CP-Retriever-Prob, slightly underperforms CP-Retriever-Code on code-related tasks, it shows marked improvements in tasks involving problem-level semantic similarity.
A.2 Ablation study
A.2.1 Ablation study on Group-infoNCE
Table 4: Performance comparison under different loss functions in contrastive learning.
Loss m (Positives) Data Aug Text-to-Code Code-to-Code
Base / / 39.60 68.05
InfoNCE 1 68.64 67.58
InfoNCE 1 68.84 68.97
InfoNCE (multi-pos) 16 66.44 66.62
Group-InfoNCE 4 69.00 69.18
Group-InfoNCE 4 70.11 70.42
Group-InfoNCE 16 69.47 69.77
Group-InfoNCE 16 70.40 70.59
Table 4 presents the performance of different loss functions and training settings on the Text-to-Code and Code-to-Code retrieval tasks. Here, InfoNCE corresponds to Equation 1, InfoNCE (multi-pos) refers to the multi-positive version in Equation 2, and Group-InfoNCE is our proposed method defined in Equation 3.
We make the following key observations:
  • Data augmentation consistently improves performance across comparable configurations, demonstrating its effectiveness.
  • InfoNCE (multi-pos) underperforms compared to the standard single-positive InfoNCE, suggesting that naïvely adding multiple positives may not be beneficial.
  • Increasing the number of positives $m$ generally leads to better performance, indicating that models benefit from more positive examples.
The ablation results confirm that our proposed Group-InfoNCE loss leads to substantial improvements in both Text-to-Code and Code-to-Code tasks, validating its effectiveness for joint representation learning of problems and solutions.
A.2.2 Ablation Study on Fine-Tuning Data Composition
Table 5: Ablation results on contrastive pretraining and fine-tuning data composition. PCD indicates whether the model was pretrained on CPRet-PCPCD using contrastive learning. Dup and Simp indicate the inclusion of Problem-to-Duplicate and Simplified-to-Full data during fine-tuning, respectively. The fourth column (PCD) denotes whether CPRet-PCPCD data was also included in the fine-tuning phase. T2C = Text-to-Code, C2C = Code-to-Code, P2Dup = Problem-to-Duplicate, S2Full = Simplified-to-Full.
T-1 T-2
PCD Dup Simp PCD T2C C2C P2Dup S2Full Avg
39.60 68.05 45.26 86.43 59.84
27.25 64.48 53.04 85.78 57.64
70.40 70.59 38.68 81.45 65.28
56.50 70.68 60.06 90.74 69.50
60.90 70.46 57.84 89.13 69.58
68.05 72.77 57.01 92.65 72.62
71.15 71.00 41.77 87.75 67.92
70.58 70.76 50.81 88.27 70.11
Ablation Study Overview. Table 5 summarizes a comprehensive ablation study on two key factors in our training pipeline: (I) contrastive pretraining on CPRet-PCPCD, and (II) the composition of fine-tuning data for problem-level tasks.
Effectiveness of Contrastive Pretraining. Rows 3–8 show that models trained with contrastive pretraining significantly outperform those without it (rows 1–2) across all four tasks. Although pretraining alone underperforms on problem-level tasks like P2Dup and S2Full, models fine-tuned from this initialization consistently surpass those trained from scratch. This indicates that contrastive pretraining helps learn transferable representations. The initial gap on problem-level tasks likely stems from objective mismatch—Group-InfoNCE focuses on problem-code alignment rather than direct problem-to-problem similarity—but still provides a strong foundation for downstream fine-tuning.
Impact of Fine-Tuning Data Composition. Rows 4–8 explore different combinations of Problem-to-Duplicate, Simplified-to-Full, and optionally CPRet-PCPCD data during fine-tuning. Using only Simplified-to-Full (row 6) yields the best overall performance, improving C2C and S2Full, while maintaining strong results on P2Dup. In contrast, adding P2Dup data (row 4) improves that task but slightly reduces generalization elsewhere. Including CPRet-PCPCD during fine-tuning (row 7, 8) helps recover performance on code-centric tasks but slightly weakens results on problem-level tasks.
Conclusion. These results confirm the value of task-specific optimization. We therefore release two final models:
  • CPRetriever-Code (row 3): trained only with contrastive pretraining, optimized for code-centric tasks.
  • CPRetriever-Prob (row 4): further fine-tuned on problem-level tasks, optimized for problem retrieval.
A.3 Further Details on the CPRet-PCPCD
A.3.1 Language and Problem Format Statistics
Figure 5(a) shows that English is the most prevalent language, largely due to the influence of prior datasets like TACO, while Chinese and Japanese examples mainly come from LibreOJ/Nowcoder and AtCoder. Figure 5(b) reveals that most problems adopt the ICPC-style full-score format, but a notable subset includes OI-style problems that offer partial scores through sub-tasks or test case breakdowns.
(a) Languages used in problem statements.
(b) ICPC-style vs OI-style grading.
Figure 5: Dataset-level characteristics by language and format in CPRet-PCPCD.
A.3.2 Programming Language Distribution
Figure 6: Distribution of programming languages in our dataset CPRet-PCPCD (log-scaled). Python is dominant due to its prevalence in prior datasets like APPS, CodeContests, and TACO. In contrast, our newly collected data is mainly in C/C++, which reflects the practical language choices in competitive programming.
As shown in Figure 6, our dataset covers a wide range of programming languages. Early datasets such as APPS, CodeContests, and TACO focused primarily on Python, driven by the popularity of Python in code generation research and the convenience it offers for model training. However, in our newly collected dataset, the dominant languages shift toward C and C++, which are widely favored in actual competitive programming due to their execution efficiency and low-level control.
Other languages such as Java, Perl, Groovy, and Rust also appear in notable quantities, indicating broader diversity and coverage in our collection.
A.4 Data Source Details
Data Copyright Notice. All data collected in this work are publicly accessible and do not involve any private or user-specific information. However, the copyright of the original content remains with the respective platforms from which the data were obtained. We use these data solely for research purposes, in accordance with the terms of use and fair academic practice.
A.4.1 Competitive Programming Problems and Codes Data(CPRet-PCPCD)
Existing Sources.
Newly Collected Sources. To enrich the dataset, we additionally collect problems and code from the following platforms:
For these new sources, we typically first obtain problem and submission metadata via the platform’s API (when available), and then download actual code submissions accordingly. In cases where APIs are not provided, we resort to web crawling.
Crawler Tools. We utilize the following libraries to implement our crawlers:
To minimize load on the original platforms and avoid abuse, we do not publicly release our custom crawlers. For academic collaboration or access, please contact us via email.
A.4.2 Duplicate Programming Problem Pairs
To collect semantically similar or duplicate programming problem pairs, we crawled discussion forum data from the following two competitive programming platforms:
From Codeforces, we gathered approximately 130,000 blog posts, each containing an average of 12 user comments. From Luogu, we collected around 236,000 discussion comments.
To extract high-quality duplicate problem pairs, we use a multi-stage filtering process combining:
  • Keyword filtering to identify relevant discussions,
  • Large Language Model (LLM)-based scoring for candidate selection, and
  • Manual annotation to ensure correctness and semantic equivalence.
Note. For each problem pair mentioned in the forums, both problems may originate from the same platform, or from different platforms. In most cases, at least one problem in the pair comes from the platform where the discussion was found.
A.4.3 Simplified and Full Problem Description Pairs
This dataset is collected entirely from the Luogu platform (https://www.luogu.com.cn/), where users have contributed simplified versions of competitive programming problems originally published on international online judges such as Codeforces (https://codeforces.com/), AtCoder (https://atcoder.jp/), SPOJ (https://www.spoj.com/), and UVA (https://onlinejudge.org/). These problems may have been originally written in English, Japanese, or other languages. The simplified versions are often written in Chinese and aim not only to translate the content, but also to restructure and clarify problem statements by removing less essential narrative elements, rephrasing complex sentences, and highlighting key constraints and requirements. This simplification process helps make the problems more accessible to beginners.
We crawled these user-generated simplifications and applied filtering steps to remove low-quality entries, such as those that are incomplete, inconsistent with the original, or overly literal machine translations. To ensure consistency across the dataset, we use Qwen-2.5-Max[29] to translate Chinese simplified statements into English before inclusion. The resulting dataset maintains a high level of clarity and fidelity, and serves as a valuable resource for studying cross-lingual and cross-abstraction retrieval.
A.5 Extended Results: Model Performance across Similarity Bins
Figure 7 expands on the similarity-based analysis in Section 3, showing average pass rates across max similarity bins for additional models not included in the main paper.
The observed trends further support the conclusion that higher retrieval similarity strongly correlates with code generation success, demonstrating consistency across different model scales and types.
Figure 7: Extended results on model performance vs. similarity.
A.6 Cases of Retrieval Tasks
A.6.1 Text-to-Code and Code-to-Code
We next present a problem along with two correct solutions. In the Text-to-Code task, the goal is to retrieve these solutions given the problem description. In the Code-to-Code task, the goal is to retrieve the other solution given one of them, which may differ in language or implementation.
https://atcoder.jp/contests/arc162/tasks/arc162_a Problem Statement There are $N$ people, numbered from $1$ to $N$, who participated in a round-trip race between two points. The following information is recorded about this race. The outward times of any two people were different, and person $i$ $(1\leq i\leq N)$ had the $i$-th fastest outward time. The round-trip times (the sum of the outward and return times) of any two people were different, and person $i$ $(1\leq i\leq N)$ had the $P_{i}$-th fastest round-trip time. The person (or persons) with the fastest return time was awarded the fastest return award. Here, $P_{1},P_{2},\dots,P_{N}$ is a permutation of $1,2,\dots,N$. How many people could have received the fastest return award? There are $T$ test cases. Answer each of them. Constraints $1\leq T\leq 500$ $2\leq N\leq 10^{3}$ $P_{1},P_{2},\dots,P_{N}$ is a permutation of $1,2,\dots,N$ All input values are integers The sum of $N$ over all test cases does not exceed $10^{3}$ Input Format The input is given from standard input in the following format:
T
case_1
case_T
Each test case $i$ $(1\leq i\leq T)$ is given in the following format:
N
P_1 P_2 ... P_N
Output Format Print $T$ lines. The $i$-th line $(1\leq i\leq T)$ should contain the answer for the $i$-th test case.
Listing 1: Solution Code 1(C++)
#include <bits/stdc++.h>
#include <cstdlib>
#include <math.h>
using namespace std;
int main()
int t;
cin >> t;
while(t--){
int n;
cin >> n;
vector<int>a(n);
vector<int>b(n);
for(int i=0;i<n;i++){
cin >> a[i];
a[i]--;
b[a[i]]=i;
int ans=0;
for(int i=0;i<n;i++){
bool ok=true;
for(int j=0;j<i;j++){
if(b[i]<b[j])ok=false;
if(ok)ans++;
cout << ans << endl;
return 0;
Listing 2: Solution Code 2(Ruby)
t = read_line.to_i
def solve
n = read_line.to_i
p = read_line.split.map { |x| x.to_i - 1 }
n.times.count do |i|
n.times.all? do |j|
# i < j => p[i] < p[j]
!(i < j) || p[i] < p[j]
end
end
end
t.times do
puts solve
end
A.6.2 Problem-to-Duplicate
The only difference between the two problems is that one includes multiple test cases, while the other does not. Aside from this, the problems are nearly identical, and are therefore classified as a Near Match.
https://codeforces.com/contest/1702/problem/G2 Problem: G2. Passable Paths (hard version) You are given a tree with $n$ vertices numbered from $1$ to $n$. A tree is an undirected connected graph with $n-1$ edges and no cycles. You are also given $q$ queries. In each query, you are given a set of $k$ vertices. For each query, determine whether there exists a simple path in the tree that passes through all the given $k$ vertices. Input The first line contains an integer $n$ ($2\leq n\leq 2\times 10^{5}$) — the number of vertices in the tree. Each of the next $n-1$ lines contains two integers $u$ and $v$ ($1\leq u,v\leq n$) — the endpoints of an edge in the tree. The next line contains an integer $q$ ($1\leq q\leq 2\times 10^{5}$) — the number of queries. Each of the next $q$ lines describes a query: The first integer $k$ ($2\leq k\leq n$) — the number of vertices in the query. Followed by $k$ integers $v_{1},v_{2},\ldots,v_{k}$ ($1\leq v_{i}\leq n$) — the vertices in the query. Output For each query, print YES if there exists a simple path that passes through all the given $k$ vertices. Otherwise, print NO.
https://www.codechef.com/JULY21A/problems/KPATHQRY Problem: KPATHQRY — Path Queries on Trees You’re given a tree with $N$ vertices numbered from $1$ to $N$. Your goal is to handle $Q$ queries. For each query, you are given $K$ nodes $v_{1},v_{2},\ldots,v_{K}$. Find whether there exists a simple path in the tree that covers all the given vertices. Input The first line contains a single integer $T$ — the number of test cases. For each test case: The first line contains a single integer $N$ — the number of vertices. Each of the following $N-1$ lines contains two integers $u$ and $v$ — an edge in the tree. The next line contains a single integer $Q$ — the number of queries. Each of the following $Q$ lines describes a query: * Starts with $K_{i}$ — the number of vertices in the query. * Followed by $K_{i}$ integers: $v_{1},v_{2},\ldots,v_{K_{i}}$. Output For each query, print "YES" if a simple path covering all the given nodes exists, otherwise print "NO". You may print each character of the string in uppercase or lowercase (for example, "yEs", "yes", "Yes", and "YES" will all be treated as identical). Constraints $1\leq T\leq 10$ $1\leq N\leq 10^{5}$ $1\leq u,v,v_{j}\leq N$ $1\leq Q\leq 10^{5}$ $1\leq K_{i}\leq N$ The edges form a valid tree. All vertices in a single query are distinct. The sum of $N$ over all test cases does not exceed $2\cdot 10^{5}$. The sum of $K_{i}$ over all queries in a single test case does not exceed $10^{5}$. Subtasks Subtask #1 (100 points): Original constraints.
A.6.3 Simplified-to-Full
The first paragraph below is a simplified version of the second paragraph. The simplification removes irrelevant background information and input/output format details, preserving only the core problem statement.
https://codeforces.com/problemset/problem/1183/F Concise problem description You have $n$ numbers from $a_{1}$ to $a_{n}$, and you want to select at most 3 numbers such that no number is a multiple of another. You aim to maximize the sum of the selected numbers. Output this maximum sum.
https://codeforces.com/problemset/problem/1183/F Full problem description Topforces Strikes Back Problem Description One important contest will take place on the most famous programming platform (Topforces) very soon! The authors have a pool of $n$ problems and should choose at most three of them into this contest. The prettiness of the $i$th problem is $a_{i}$ The authors have to compose the most pretty contest (in other words, the cumulative prettinesses of chosen problems should be maximum possible). But there is one important thing in the contest preparation: because of some superstitions of authors, the prettinesses of problems cannot divide each other. In other words, if the prettinesses of chosen problems are $x,y,z$ , then $x$ should be divisible by neither $y$ , nor $z$ , $y$ should be divisible by neither $x$ , nor $z$ and $z$ should be divisible by neither $x$ , nor $y$ If the prettinesses of chosen problems are $x$ and $y$ then neither $x$ should be divisible by $y$ nor $y$ should be divisible by $x$ Any contest composed from one problem is considered good. Your task is to find out the maximum possible total prettiness of the contest composed of at most three problems from the given pool. You have to answer $q$ independent queries. If you are Python programmer, consider using PyPy instead of Python when you submit your code. Input Format The first line of the input contains one integer $q$ ( $1\leq q\leq 2\cdot 10^{5}$ ) — the number of queries. The first line of the query contains one integer $n$ ( $1\leq n\leq 2\cdot 10^{5}$ ) — the number of problems. The second line of the query contains $n$ integers $a_{1},a_{2},\dots,a_{n}$ ( $2\leq a_{i}\leq 2\cdot 10^{5}$ ), where $a_{i}$ is the prettiness of the $i$th problem. It is guaranteed that the sum of $n$ over all queries does not exceed $2\cdot 10^{5}$ Output Format For each query print one integer — the maximum possible cumulative prettiness of the contest composed of at most three problems from the given pool of problems in the query.